OpenStudy (anonymous):

Simplify as much as possible: (a/b)^4

jimthompson5910 (jim_thompson5910):

you are raising a/b to the 4th power that is the same as raising a to the 4th power over b to the 4th power ie \[\Large \left(\frac{a}{b}\right)^4 = \frac{a^4}{b^4}\] the general rule is \[\Large \left(\frac{a}{b}\right)^k = \frac{a^k}{b^k}\]